# Rehearsal-Free Continual Federated Learning with Synergistic Regularization

### v1

@@ 1 Introduction
Federated learning (FL) is to facilitate the collaborative training of a global deep learning model among multiple edge clients while ensuring the privacy of their locally stored data (McMahan et al., 2017; Wang et al., 2023a; Liu et al., 2024). Recently, FL has garnered significant interest and found applications in diverse domains, including recommendation systems (Yang et al., 2020; Li et al., 2024d) and smart healthcare solutions (Xu et al., 2021; Nguyen et al., 2022).
@@
To address this issue, researchers have studied continual federated learning (CFL), which enables each client to continuously learn from local streaming tasks. FedCIL is proposed in (Qi et al., 2023) to learn a generative network and reconstruct previous samples for replay, improving the retention of previous information. The authors in (Li et al., 2024a) propose to selectively cache important samples from each task and train the local model with both cached samples and the samples from the current task. The authors in (Dong et al., 2022; 2023a) focus on the federated class-incremental learning scenario and train a global model by computing additional class-imbalance losses. The study in (Ma et al., 2022a) employs supplementary distilled data on both server and client ends, leveraging knowledge distillation to mitigate catastrophic forgetting.
@@
In this paper, we aim to enhance the regularization technique for continual federated learning by adhering to the following three principles: (1) free rehearsal, (2) low computational cost, and (3) robustness to data heterogeneity. In response to the first two principles mentioned above, we observe that a considerable number of traditional regularization techniques in the centralized environment can offer a possible solution intuitively. To verify it, we deploy these techniques into the CFL scenario and conduct extensive experiments, which proves that some of the traditional regularization techniques can indeed achieve fairish results. Despite the advantages of these techniques, we find that directly combining it with CFL fails to maintain stable performance under varying data heterogeneity. For instance, owing to the traditional synaptic intelligence algorithm, each client can prevent catastrophic forgetting in a free rehearsal and low-cost manner by calculating the surrogate loss of different incremental tasks. However, this surrogate loss is merely calculated based on the local data distribution, and in the case of highly heterogeneous data, the local optimization objective of the surrogate loss may not align with that of the global data distribution. It is quite necessary and important to consider and solve the problem of data heterogeneity when deploying algorithms in real-world CFL scenarios.
To address this issue, we propose an enhanced Synaptic Intelligence (SI)-based CFL approach with synergistic regularization named FedSSI. Specifically, in FedSSI, each client will calculate the surrogate loss for previous tasks based not only on information from the local dataset but also on its correlation to the global dataset. We introduce a personalized surrogate model (PSM) for each client, which incorporates global knowledge into local caching so that the surrogate loss reflects both local and global understandings of the data. The SI algorithm is then employed to mitigate catastrophic forgetting, with each client training the local model using both the new task loss and the PSM. Through extensive experiments on various datasets and two types of incremental tasks (Class-IL and Domain-IL), we demonstrate that FedSSI significantly improves model accuracy compared to state-of-the-art approaches. The major contributions of this paper are summarized as follows:
  • We provide an in-depth analysis of regularization-based CFL, identifying that existing methods primarily rely on data rehearsal or massive computational overhead, which poses significant challenges where resources are constrained and data privacy must be maintained.
  • We improve regularization techniques to address catastrophic forgetting and data heterogeneity in CFL. We propose FedSSI, a simple and efficient method that regulates model updates with both local and global information about data heterogeneity.
@@ 2 Background and Related Work
Federated Learning. Federated Learning (FL) is a technique designed to train a shared global model by aggregating models from multiple clients that are trained on their own local private datasets (McMahan et al., 2017; Wang et al., 2023b; 2024b). One widely used architecture for FL is FedAvg (McMahan et al., 2017), which optimizes the global model by aggregating the parameters of local models trained on private local data. However, traditional FL algorithms like FedAvg face challenges due to data heterogeneity, where the datasets on clients are Non-IID (non-independent and identically distributed), resulting in degraded model performance (Jeong et al., 2018b; Wang et al., 2024a). To address the Non-IID issue in FL, a proximal term is introduced in the optimization process in (Li et al., 2020) to mitigate the effects of heterogeneous and Non-IID data distribution across participating devices. Another approach, federated distillation (Jeong et al., 2018a), aims to distill the knowledge from multiple local models into the global model by aggregating only the soft predictions generated by each model. The authors in (Lin et al., 2020) proposed a knowledge distillation method that utilizes unlabeled training samples as a proxy dataset. However, these methods primarily address Non-IID static data with spatial heterogeneity, overlooking potential challenges posed by streaming tasks with temporal heterogeneity.
@@
Continual Learning. Continual Learning (CL) is a machine learning technique that allows a model to continuously learn from streaming tasks while retaining knowledge gained from previous tasks (Hsu et al., 2018; van de Ven & Tolias, 2019). This includes task-incremental learning (Dantam et al., 2016; Maltoni & Lomonaco, 2018), class-incremental learning (Rebuffi et al., 2017; Yu et al., 2020), and domain-incremental learning (Mirza et al., 2022; Churamani et al., 2021). Existing approaches in CL can be classified into three main categories: replay-based methods (Rebuffi et al., 2017; Liu et al., 2020), regularization-based methods (Jung et al., 2020; Yin et al., 2020), and parameter isolation methods (Long et al., 2015; Fernando et al., 2017). Replay-based methods select representative old samples to retain previously learned knowledge when training on a new task. Regularization-based methods protect existing knowledge from being overwritten by new knowledge by imposing constraints on the loss function of new tasks. Parameter isolation methods typically introduce additional parameters and computations to learn new tasks. Our focus is on the continual federated learning scenario, which combines the principles of federated learning and continual learning.
@@
Continual Federated Learning. Continual Federated Learning (CFL) aims to address the learning of streaming tasks in each client by emphasizing the adaptation of the global model to new data while retaining knowledge from past data. Despite its significance, CFL has only recently garnered attention, with (Yoon et al., 2021) being a pioneering work in this field. Their research focuses on Task-IL, requiring unique task IDs during inference and utilizing separate task-specific masks to enhance personalized performance. It is studied in (Bakman et al., 2023) that projecting the parameters of different tasks onto different orthogonal subspaces prevents new tasks from overwriting previous task parameters. Other studies, such as (Ma et al., 2022b), utilize knowledge distillation at the server and client levels using a surrogate dataset. Recently, (Li et al., 2024b; a) proposed calculating the importance of samples separately for the local and global distributions, selectively saving important samples for retraining to mitigate catastrophic forgetting in federated incremental learning scenarios. Some research, like (Jiang et al., 2021; Dong et al., 2023b), explores CFL in domains beyond image classification. The authors in (Li et al., 2024c) adopt the concept of dynamic networks to allow each client to train multiple personalized models based on local resource availability, effectively isolating knowledge between different incremental tasks and merging models for similar knowledge tasks. Most existing works need to cache extra samples with a memory buffer, and methods like (Bakman et al., 2023) and (Yoon et al., 2021) require the task ID during the inference stage. Our work focuses on regularization-based CFL methods with free rehearsal and low computational cost.
3 Problem Formulation and Preliminaries
@@
$\begin{split}\mathcal{L}_{total}(w_{k}^{t})=\mathcal{L}_{new}(w_{k}^{t})+\alpha\mathcal{L}_{sur}(w_{k}^{t})=\mathcal{L}_{new}(w_{k}^{t})+\alpha\sum_{i}\Omega_{k,i}^{t}||w_{k,i}^{t}-w_{i}^{t-1}||^{2}.\end{split}$ (2)
@@
where $t^{l-1}$ and $t^{l}$ denote the start and end iteration of the $l$-th task. By summing the absolute values of the gradients over all training iterations, SI estimates the overall importance of each weight for previous tasks. This importance measure is then used to penalize significant changes to those weights during training on new tasks, thus reducing catastrophic forgetting.
4 Regularization Techniques for continual federated learning
@@
Regularization techniques have been known to be cost-efficient for CL without data rehearsal. We in this section analyze several typical regularization techniques used in centralized continual learning and then examine their application in continual federated learning (CFL) with different data heterogeneity settings. Through these examinations, we identify an interesting observation: existing regularization techniques (especially synaptic intelligence) exhibit great advantages in IID settings but fail to hinder knowledge forgetting in Non-IID settings. Based on these observations, we then propose a novel method that tailors the synaptic intelligence for the CFL scenario, namely, FedSSI.
4.1 Deep Dive Into Traditional Regularization Techniques in CFL
@@
Based on three principles mentioned in Section 1, here we use four notable algorithms: LwF (Li & Hoiem, 2017), EWC (Kirkpatrick et al., 2017), OGD (Farajtabar et al., 2019), and SI (Zenke et al., 2017) to explore the performance of traditional centralized regularization techniques in the CFL scenario. We simply apply them during the local training process of each client. Each algorithm is combined with the FedAvg (McMahan et al., 2017) framework, and we report the final accuracy $A(f)$ after the model completes training on the last streaming task across all tasks. More details about experimental settings can be found in Section 5.1.
Figure 1: Performance comparisons of various regularization-based CFL methods on CIFAR10 and Digit10 datasets with IID data.
@@
Regularization can mitigate forgetting in decentralized scenarios when data is IID among clients. As shown in Figure 2, all regularization methods significantly improve test accuracy compared to the baseline, where FL+EWC, FL+OGD, and FL+SI improve the performance by approximately 6$\sim$10%. This suggests that regularization techniques in distributed settings with multiple clients are still effective, indicating their robustness to the restriction of data isolation. Moreover, it is worth noting that FL with SI performs the best among all regularization methods, indicating a great potential to tackle the knowledge-forgetting problem in CFL.
Regularization fails to mitigate knowledge forgetting when data is Non-IID among clients. As shown in Figure 2, we observe that as the level of data heterogeneity increases, the performance of all methods significantly declines. Their performance is almost the same as the baseline FedAvg when the data heterogeneity $\alpha=0.1$, which is a common setting in FL. This phenomenon also occurs in FL+SI, which completely loses its advantage in IID settings. Therefore, to fully unleash the potential of existing regularization methods in CFL, the data heterogeneity should be taken into account when combining them.
@@ 4.2 FedSSI: Improved Regularization Technique with synergism for CFL
Input : $T$: communication round;  $K$: client number;  $\eta$: learning rate;  $\{\mathcal{T}^{t}\}_{t=1}^{n}$: distributed dataset with $n$ tasks;  $w$: parameter of the model;  $v_{k}^{t}$: personalized surrogate model in client $k$ for the $t$-th task;  $s^{l}_{k,i}$: contribution of the $i$-th parameter in client $k$ with $t$-th task.
Output : ${w_{1},w_{2},\ldots,w_{k}}$: target classification model for each client.
1 for $c=1$ to $T$ do // Before the arrival of the $t$-th new task
2         Server randomly selects a subset of devices $S_{t}$ and send $w^{t-1}$
3       for each selected client $k\in S_{t}$ in parallel do
4             $\text{Update}\ v_{k}^{t-1}\text{in $s$ local iterations}$ with (5).
5             for During the update of $v_{k}^{t-1}$ do
6                   Calculate the contribution for each parameter $s^{l}_{k,i}$ with the ($t$-1)-th task by (6).
7
8             end for
9            Training the target classification model $w_{k}^{t}$ with the new task with (2)(1). // Receive the $t$-th new task
10             Send the model $w_{k}^{t}$ back to the server.
11       end for
12      $w^{t}\leftarrow\text{ServerAggregation}(\{w_{k}^{t}\}_{k\in S_{t}})$
13 end for
@@ 12      $w^{t}\leftarrow\text{ServerAggregation}(\{w_{k}^{t}\}_{k\in S_{t}})$
$v_{k,s}^{t-1}=v_{k,s-1}^{t-1}-\eta\biggl{(}\sum_{i=1}^{M}\nabla\mathcal{L}\left(f_{v_{k,s-1}^{t-1}}({x}_{k,t-1}^{(i)}),{y}_{k,t-1}^{(i)}\right)+q(\lambda)(v_{k,s-1}^{t-1}-w^{t-1})\biggl{)}.$ (5)
where $\ q(\lambda)=\frac{1-\lambda}{2\lambda},\lambda\in(0,1)$, and $\eta$ is the rate to control the step size of the update. The hyper-parameter $\lambda$ adjusts the balance between the local and global information in the update.
@@ 4.3 Analytical Understanding of the Personalized Surrogate Model in FedSSI
$\begin{split}\hat{v_{k}}(\lambda):=\mathop{\arg\min}\limits_{v_{k}}\Big{\{}f(v_{k})+\frac{q(\lambda)}{2}||v_{k}-\hat{w}||^{2}\Big{\}},\quad\text{where}\ q(\lambda):=\frac{1-\lambda}{2\lambda},\ \lambda\in(0,1).\end{split}$ (7)
@@
Proof. The proof here directly follows the proof in Theorem 3.1 (Hanzely & Richtárik, 2020). As $\lambda$ declines and $q(\lambda)$ grows, the objective of Eq.  7 tends to optimize $||v_{k}-\hat{w}||^{2}$ and increase the local empirical training loss $f(v_{k})$, leading to the convergence on the global model. Hence, we can modify the $\lambda$ value to adjust the optimization direction of our model $v_{k}$, thus the dominance of local and global model information.
@@
5 Experiments
5.1 Experiment Setup
Configurations. Unless otherwise mentioned, we employ ResNet18 (He et al., 2016) as the backbone model in all methods and use the Dirichlet distribution Dir($\alpha$) to distribute local samples, inducing data heterogeneity for all tasks, where a smaller $\alpha$ indicates higher data heterogeneity. For a fair comparison, we set the memory buffer $300$ for each client in models designed for continual federated learning with data rehearsal to cache synthetic or previous samples. We report the final accuracy $A(f)$ upon completion of the last streaming task and the average accuracy $\bar{A}$ across all tasks. Each experiment set is run twice, and we take each run’s final 10 rounds’ accuracy to calculate the average value and standard variance. We use Adam as the optimizer with a linear learning rate schedule. All experiments are run on 8 RTX 4090 GPUs and 16 RTX 3090 GPUs. Detailed settings and benchmark parameters are illustrated in Table 6. (In Appendix C)
@@ 5.1 Experiment Setup
5.2 Performance Overview
Test Accuracy. Table 2 shows the test accuracy of various methods with data heterogeneity across six datasets. We report both the final accuracy and average accuracy of the global model when all clients finish their training on all tasks. FOT outperforms other baselines on three class-IL datasets as the orthogonal projection in FOT exhibits effective training with obvious boundaries between streaming tasks. However, its performance experiences a significant decline with domain-IL datasets, where the classes of each task remain unchanged. All regularization-based methods work and CFL models with data rehearsal fail to achieve ideal results with limited memory buffer. FedSSI achieves the best performance in all cases by up to 11.52% in terms of final accuracy. More discussions and results on model performance and communication efficiency are available in Appendix E.

### v2

# FedSSI: Rehearsal-Free Continual Federated Learning with Synergistic Synaptic Intelligence

@@ 1 Introduction
Federated learning (FL) is to facilitate the collaborative training of a global deep learning model among multiple edge clients while ensuring the privacy of their locally stored data (McMahan et al., 2017; Wang et al., 2023a; Liu et al., 2024). Recently, FL has garnered significant interest and found applications in diverse domains, including recommendation systems (Yang et al., 2020; Li et al., 2024c) and smart healthcare solutions (Xu et al., 2021; Nguyen et al., 2022).
@@
To address this issue, researchers have studied continual federated learning (CFL), which enables each client to continuously learn from local streaming tasks. FedCIL is proposed in (Qi et al., 2023) to learn a generative network and reconstruct previous samples for replay, improving the retention of previous information. The authors in (Li et al., 2024a) propose to selectively cache important samples from each task and train the local model with both cached samples and the samples from the current task. The authors in (Dong et al., 2022, 2023) focus on the federated class-incremental learning scenario and train a global model by computing additional class-imbalance losses. The study in (Ma et al., 2022a) employs supplementary distilled data on both server and client ends, leveraging knowledge distillation to mitigate catastrophic forgetting.
@@
In this paper, we aim to enhance the regularization technique for continual federated learning by adhering to the following three principles: (1) free rehearsal, (2) low computational cost, and (3) robustness to data heterogeneity. In response to the first two principles mentioned above, we observe that a considerable number of traditional regularization techniques in the centralized environment can offer a possible solution intuitively. To verify it, we deploy these techniques into the CFL scenario and conduct extensive experiments, which proves that some of the traditional regularization techniques can indeed achieve fairish results, where Synaptic Intelligence (SI) (Zenke et al., 2017) is particularly outstanding and greatly outperforms other methods. Despite the advantages of these techniques, we find that directly combining them with CFL fails to maintain stable performance under varying data heterogeneity, with SI experiencing the most significant decline. In principle, this is due to surrogate loss, as the main component of SI, being merely calculated based on the local data distribution, causing the local optimization objective to not align with that of the global data distribution in the case of highly heterogeneous data.
To address this issue, we propose the Synergistic Synaptic Intelligence for the continual Federated Learning (FedSSI). Specifically, in FedSSI, each client will calculate the surrogate loss for previous tasks based not only on information from the local dataset but also on its correlation to the global dataset. We introduce a personalized surrogate model (PSM) for each client, which incorporates global knowledge into local caching so that the surrogate loss reflects both local and global understandings of the data. The SI algorithm is then employed to mitigate catastrophic forgetting, with each client training the local model using both the new task loss and the PSM. Through extensive experiments on various datasets and two types of incremental tasks (Class-IL and Domain-IL), we demonstrate that FedSSI significantly improves model accuracy compared to state-of-the-art approaches. Our major contributions are:
  • We provide an in-depth analysis of regularization-based CFL, identifying that despite existing methods, especially SI, being effective on homogeneous data, they cannot guarantee performance on heterogeneous data distributions which are common in FL.
  • We propose FedSSI, a simple yet effective method that tailors SI for the CFL. Specifically, FedSSI regulates model updates with both local and global information about data heterogeneity.
@@ 2 Background and Related Work
Federated Learning. Federated Learning (FL) is a technique designed to train a shared global model by aggregating models from multiple clients that are trained on their own local private datasets (McMahan et al., 2017; Wang et al., 2023b, 2024b). However, traditional FL algorithms like FedAvg face challenges due to data heterogeneity, where the datasets on clients are Non-IID, resulting in degraded model performance (Jeong et al., 2018b; Wang et al., 2024a). To address the Non-IID issue in FL, a proximal term is introduced in the optimization process in (Li et al., 2020) to mitigate the effects of heterogeneous and Non-IID data distribution across participating devices. Another approach, federated distillation (Jeong et al., 2018a), aims to distill the knowledge from multiple local models into the global model by aggregating only the soft predictions. However, these methods primarily address Non-IID static data with spatial heterogeneity, overlooking potential challenges posed by streaming tasks with temporal heterogeneity.
@@
Continual Learning. Continual Learning (CL) is a machine learning technique that allows a model to continuously learn from streaming tasks while retaining knowledge gained from previous tasks (Hsu et al., 2018; van de Ven & Tolias, 2019). This includes task-incremental learning (Dantam et al., 2016; Maltoni & Lomonaco, 2018), class-incremental learning (Rebuffi et al., 2017; Yu et al., 2020), and domain-incremental learning (Mirza et al., 2022; Churamani et al., 2021). Existing approaches in CL can be classified into three main categories: replay-based methods (Rebuffi et al., 2017; Liu et al., 2020), regularization-based methods (Jung et al., 2020; Yin et al., 2020), and parameter isolation methods (Long et al., 2015; Fernando et al., 2017). Our focus is on the continual federated learning scenario, which combines the principles of federated learning and continual learning.
@@
Continual Federated Learning. Continual Federated Learning (CFL) aims to address the learning of streaming tasks in each client by emphasizing the adaptation of the global model to new data while retaining knowledge from past data. Despite its significance, CFL has only recently garnered attention, with (Yoon et al., 2021) being a pioneering work in this field. Their research focuses on Task-IL, requiring unique task IDs during inference and utilizing separate task-specific masks to enhance personalized performance. It is studied in (Bakman et al., 2023) that projecting the parameters of different tasks onto different orthogonal subspaces prevents new tasks from overwriting previous knowledge. Other studies, such as (Ma et al., 2022b), utilize knowledge distillation at the server and client levels using a surrogate dataset. Recently, (Li et al., 2024b, a) proposed calculating the importance of samples separately for the local and global distributions, selectively saving important samples for retraining to mitigate catastrophic forgetting. Most existing works need to cache extra samples with a memory buffer, and methods like (Bakman et al., 2023) and (Yoon et al., 2021) require the task ID during the inference stage. Our work focuses on regularization-based CFL methods with free rehearsal and low computational cost.
Figure 1: Performance comparisons of regularization-based CFL methods on CIFAR10 and Digit10 datasets with IID data.
3 Preliminaries
@@
$\begin{split}\mathcal{L}_{total}(w_{k}^{t})&=\mathcal{L}_{new}(w_{k}^{t})+\alpha\mathcal{L}_{sur}(w_{k}^{t})\\ &=\mathcal{L}_{new}(w_{k}^{t})+\alpha\sum_{i}\Omega_{k,i}^{t}||w_{k,i}^{t}-w_{i}^{t-1}||^{2}.\end{split}$ (2)
@@
where $t^{l-1}$ and $t^{l}$ denote the start and end iteration of the $l$-th task. By summing the absolute values of the gradients over all training iterations, SI estimates the overall importance of each weight for previous tasks. Then, it is used to penalize significant changes to those weights during training on new tasks, thus reducing catastrophic forgetting.
4 Methodology
@@
Regularization techniques have been known to be cost-efficient for CL without data rehearsal. In this section, we analyze several typical regularization techniques used in centralized continual learning and then examine their application in continual federated learning (CFL) with different data heterogeneity settings. Through these examinations, we identify an interesting observation: existing regularization techniques (especially synaptic intelligence) exhibit great advantages in IID settings but fail to hinder knowledge forgetting in Non-IID settings. Based on these observations, we then propose a novel method that tailors the synaptic intelligence for the CFL scenario, namely, FedSSI.
4.1 Observations
@@
Based on three principles mentioned in Section 1, here we use four notable algorithms: LwF (Li & Hoiem, 2017), EWC (Kirkpatrick et al., 2017), OGD (Farajtabar et al., 2019), and SI (Zenke et al., 2017) to explore the performance of traditional centralized regularization techniques in the CFL scenario. We simply apply them during the local training process of each client. Each algorithm is combined with the FedAvg (McMahan et al., 2017) framework, and we report the final accuracy $A(f)$ after the model completes training on the last streaming task across all tasks. More details about experiments can be found in Section 5.1.
@@
Observation 1. Regularization especially SI can mitigate forgetting in decentralized scenarios when data is IID among clients. As shown in Figure 2, all regularization methods significantly improve test accuracy compared to the baseline, where FL+EWC, FL+OGD, and FL+SI improve the performance by approximately 6$\sim$10%. This suggests that regularization techniques in distributed settings with multiple clients are still effective, indicating their robustness to the restriction of data isolation. Moreover, it is worth noting that FL with SI performs the best among all regularization methods, indicating a great potential to tackle the knowledge-forgetting problem in CFL.
Observation 2. Regularization fails to mitigate knowledge forgetting when data is Non-IID among clients, with SI declining the most severely. As shown in Figure 2, we observe that as the level of data heterogeneity increases, the performance of all methods significantly declines. Their performance is almost the same as the baseline FedAvg when the data heterogeneity $\alpha=0.1$, which is a common setting in FL. This phenomenon is more pronounced in FL+SI, where SI experiences the greatest performance decline among all methods, sometimes even performing worse than other methods (e.g., $\alpha=1.0$ on the Digit10 dataset). Based on the above observations, if the data heterogeneity issues encountered by SI in FL can be resolved, then the problem of catastrophic forgetting in CFL could be significantly alleviated.
@@ 4.2 FedSSI: Improved Regularization Technique with Synergism for CFL
0:  $T$: communication round;  $K$: client number;  $\eta$: learning rate;  $\{\mathcal{T}^{t}\}_{t=1}^{n}$: distributed dataset with $n$ tasks;  $w$: parameter of the model;  $v_{k}^{t}$: personalized surrogate model in client $k$ for the $t$-th task;  $s^{l}_{k,i}$: contribution of the $i$-th parameter in client $k$ with $t$-th task.
0:  ${w_{1},w_{2},\ldots,w_{k}}$: target classification model.
1:  Initialize the parameter $w$;
2:  Distribute $w$ to all clients;
3:  for $c=1$ to $T$ do
4:     Server randomly selects a subset of devices $S_{t}$ and send $w^{t-1}$;
5:     for each selected client $k\in S_{t}$ in parallel do
6:        $\text{Update}\ v_{k}^{t-1}\text{in $s$ local iterations}$ with (5);
7:        for during the update of $v_{k}^{t-1}$ do
8:           Calculate the contribution for each parameter $s^{l}_{k,i}$ with the ($t$-1)-th task by (6);
9:        end for
10:        Training the target classification model $w_{k}^{t}$ with the new task with (2)(1);
11:        Send the model $w_{k}^{t}$ back to the server;
[…]
13:  end for
@@
$v_{k,s}^{t-1}$ $=v_{k,s-1}^{t-1}-\eta\biggl{(}\sum_{i=1}^{M}\nabla\mathcal{L}\left(f_{v_{k,s-1}^{t-1}}({x}_{k,t-1}^{(i)}),{y}_{k,t-1}^{(i)}\right)$ (5)
$+q(\lambda)(v_{k,s-1}^{t-1}-w^{t-1})\biggl{)}.$
where $\ q(\lambda)=\frac{1-\lambda}{2\lambda},\lambda\in(0,1)$, and $\eta$ is the rate to control the step size of the update. The hyper-parameter $\lambda$ adjusts the balance between the local and global information.
@@ 4.3 Analytical Understanding of the Personalized Surrogate Model in FedSSI
$\begin{split}\hat{v_{k}}(\lambda):=\mathop{\arg\min}\limits_{v_{k}}\Big{\{}f(v_{k})+\frac{q(\lambda)}{2}||v_{k}-\hat{w}||^{2}\Big{\}},\\ \quad\text{where}\ q(\lambda):=\frac{1-\lambda}{2\lambda},\ \lambda\in(0,1).\end{split}$ (7)
@@
Proof. The proof here directly follows the proof in Theorem 3.1 (Hanzely & Richtárik, 2020). As $\lambda$ declines and $q(\lambda)$ grows, the Eq.( 7) tends to optimize $||v_{k}-\hat{w}||^{2}$ and increase the local empirical training loss $f(v_{k})$, leading to the convergence on the global model. Hence, we can modify the $\lambda$ to adjust the optimization direction of our model $v_{k}$, thus the dominance of local and global information.
@@
5 Experiments
5.1 Experiment Setup
Configurations. Unless otherwise mentioned, we employ ResNet18 (He et al., 2016) as the backbone model in all methods and use the Dirichlet distribution $\alpha$ to distribute local samples, inducing data heterogeneity for all tasks, where a smaller $\alpha$ indicates higher data heterogeneity. For a fair comparison, we set the memory buffer $300$ for each client in models designed for continual federated learning with data rehearsal to cache synthetic or previous samples. We report the final accuracy $A(f)$ upon completion of the last streaming task and the average accuracy $\bar{A}$ across all tasks. All experiments are run on 8 RTX 4090 GPUs and 16 RTX 3090 GPUs. Detailed settings and benchmark parameters are illustrated in Table 5 (In Appendix C).